Convert from nominal to real and calculate economic growth for the applicable years:


2011 2012 2013 2014 2015

20K 23K 25K 28K 30K

10% 9% 10% 8%


1
Expert's answer
2021-10-14T17:06:10-0400

Solution:

Real value = Nominal value + Inflation

Real value 2011 = 20,000 + (20,000 "\times 10\%") = 20,000 + 2,000 = 22,000

Real value 2012 = 23,000 + (23,000 "\times 9\%") = 23,000 + 2,070 = 25,070

Real value 2013 = 25,000 + (25,000 "\times 10\%") = 25,000 + 2,500 = 27,500

Real value 2014 = 28,000 + (28,000 "\times 8\%") = 28,000 + 2,240 = 30,240

Real value 2015 = 30,000 + (30,000 "\times 7\%") = 30,000 + 2,100 = 32,100

 

Economic growth rate = "\frac{Current\; real\; value}{Previous\; real\; value}\times 100\%"


Economic growth rate 2012 ="\frac{25,070 - 22,000}{25,070}\times 100\% = 13.95\%"


Economic growth rate 2013 = "\frac{27,500 - 25,070}{25,070}\times 100\% = 9.69\%"


Economic growth rate 2014 = "\frac{30,240 - 27,500}{27,500}\times 100\% = 9.96\%"


Economic growth rate 2015 = "\frac{32,100 - 30,240}{30,240}\times 100\% = 6.15\%"